OpenStudy (anonymous):

What number(s) should be excluded as values of x in the polynomial? x+2/x^2-4

OpenStudy (anonymous):

\[x+2\div x^2-4\]

OpenStudy (anonymous):

so, 2x/x^2 =0 + 4= 4 so, x= 4

OpenStudy (anonymous):

x+2/x^2-4 here if we take the value of x= 2, 2+2/(2)^2-4 = 4/0 =undefined if x=-2, then -2 +2/(-2)^2-4 = 0/0=not defined so the values -2 and 2 has to be excluded while considering the value of x